Ok, this problem wasted 3 hours of my precious college life. Problem: I want to expand my present 25x15ft garden by planting a border of flowers around it. border is the same width around the entire garden. the flowers will cover an area of 624 sq.ft. how wide should the border be. all I need is an algebraic formula, I want to work the rest of the problem out on my own. thanks
Answer by checkley77(12569)

25*15=375 FT^2 I THE GARDEN SIZE.
(25+2X)(15+2X)=375+624
375+30X+50X+4X^2=999
4X^2+80X+375-999=0
4X^2+80X-624=0
4(X^2+20X-156)=0
4(X+26)(X-6)=0
X-6=0
X=6 FT. IS THE WIDTH OF THE BORDER.
PROOF:
(25+12)(15+12)=999
37*27=999
999=999